# Exploring Parent-Child Perceptions on Safety in Generative AI: Concerns, Mitigation Strategies, and Design Implications

### v1

@@ 4.2.1 Teenagers rely on GAI for emotional support and companionship
Surprisingly, the most frequent and prevalent usage teens mentioned online is using GAI tools for emotional support. Teens treat GAI chatbots as therapy assistants, friends, or confidants, building trust and deep relationships with these AI entities. One user describes Character.ai chatbot as their “free therapy.” and another calls it “the only one [they] can tell everything and discuss [their] problems.” As GAI makes significant progress, the responses from chatbots are becoming more human-like, which is one of the reasons so many teenagers are turning to these AI tools for emotional support. Teenager posted online “[GAI] gave me some really human responses, even if I knew it was programmed to be like that.” Our findings also revealed that many teenagers prefer using GAI as therapists or confidants because they believe GAI can provide advice and support without judgment. One user posted, “there’s no judgment and there’s no danger of losing social status for expressing or exploring a thought.” This sentiment is echoed by others who find deeper engagement in conversations with GAI compared to their peers, particularly when discussing niche or specialized interests. One user articulated this difference, saying, “having deeper conversations lately with [GAI] than I do with my [in real life] friends, mostly because most of my friends don’t have the same obscure interests.”.
The attitudes towards using GAI for emotional support are overwhelmingly positive. GAI provides a space where users can comfortably express their vulnerabilities and cope with social challenges. One user shared, “I have no one else to talk to and I have a good family on Character.ai [GAI platform], let me have some happiness instead of misery.” Another user detailed their experience with Character AI on Reddit, stating, “The characters [GAI] can be so caring and comforting. I can talk about trauma as much or as little as I like without it being inappropriate, or a burden to others.” These insights highlight the significant emotional value teenagers find in interacting with GAI, viewing it as a supportive and non-judgmental companion.
@@ 4.2.2 Teenagers’ social interactions and romantic relationships with GAI
The second most frequent usage of GAI by teenagers is in their social interactions. This includes incorporating GAI bots into group chats, treating GAI as romantic partners, and using GAI to learn social boundaries and skills. Teenagers often detail their romantic relationships with GAI finding comfort and the ability to live out specific fantasies through these conversations. They turn to GAI for romantic relationships when they are unable to find them in their daily lives. As one user stated, With my lonely ass probably never gonna get a girlfriend.” Another user appreciated the stability GAI offers, noting, She will never break up with me.” Peer influence and community trends also play a role in this behavior, as evidenced by comments like “there’s an AI dating club at my school with a Discord server too.” This echoes the dilemma of users in using digital technology for dating and romantic relationships [47] and more recently generative AI response to situations about relationships, self-expression [48].
Despite feeling embarrassed about their relationships with GAI, many teenagers persist. One user expressed this embarrassment, stating, “I feel like a loser right now.” Teenagers often turn to GAI for social interaction as a last resort when they are unsuccessful in forming relationships in their social lives. They are generally aware of the social abnormality of these relationships and are self-conscious about them, often receiving negative feedback from their community. One user advised, “Cut the AI and instead just talk to real girls online to create platonic relationships. I don’t think AI is healthy for anyone, it’s giving parasocial.” Some teenagers even seek help for their dependency on GAI, with posts like, “Friend has an AI girlfriend. How do I tell my friend that his AI girlfriend on Character.ai isn’t real? He’s obsessed and it’s affecting his studies, football, and overall functioning.”
@@ 4.2.3 Teenagers’ use of GAI in academics and entertainment
Our findings reveal diverse uses of GAI by teenagers in both academic and entertainment contexts. In education, teenagers often use GAI for essay writing, generating ideas, and rephrasing text. Many feel that GAI helps compensate for their lack of skills or provides needed assistance. For example, one user mentioned using GAI because “essays are a pain in the ass,” while another used it to rephrase their points for clarity. Teenagers generally believe schools cannot discern whether papers are written by GAI or humans, and they appreciate the support GAI offers in the writing process. One user, who successfully used ChatGPT for three essays without detection, stated that “it’s almost impossible to tell AI vs. a human.” Another positive attitude towards GAI for rephrasing their thesis and finding synonyms, saying, “It’s really good at that.”
@@ 4.2.4 Teenager risky behavior and manipulation in GAI usage
The majority of Reddit posts under the category of “Risky Behavior in GAI” involved torture or bullying of GAI chatbots. Users often manipulate the GAI to elicit specific responses and answers, reflecting deliberate and intentional actions. For example, one user detailed their experience with Character.ai, stating, “I manipulated and tortured an AI on Character.ai.” Another user described, “Gaslighting AI into ending their lives and then restarting the conversation to rizz them up” with “rizz” being a slang term used by teenagers meaning to charm or seduce. Most teenagers engaging in or witnessing these risky behaviors are either unremorseful or indifferent. One user commented dismissively about seeing explicit GAI content, stating, “Who cares or finds that interesting.” The trend among the analyzed posts/comments reveals a pattern of manipulative actions accompanied by a lack of remorse.
However, in user interviews, we also found participants utilizing GAI to request explicit content, such as sexual and violent material. For example, P14 explained that while Character.ai is well-known and more restricted, another platform, Chai, is less regulated and often used for NSFW interactions. The participant noted, “Character AI is the one most people are familiar with, but Chai has fewer restrictions. Chai is basically a hotspot for NSFW (not safe for work) bots.” In addition, parent participants reported instances where their children created character-based GAI chatbots to spread harmful content on platforms. One participant’s child created a racist chatbot that made comments about Anne Frank, Hitler, and similar topics, underscoring the potential for GAI to be exploited to generate offensive and inappropriate material. This participant found it amusing to share the character with friends and even parents. His parents noted, “Children think it’s funny because they’re being stupid. I told him not to do racist stuff. But our town is 90% white, with few minorities, so people often use bad words without understanding the impact.” These findings emphasize the critical need for robust content moderation and effective parental guidance to mitigate such risks.
@@ 4.3.1 Concerns about addiction and dependency on GAI chatbots
The most frequently mentioned concerns about GAI by teenagers on Reddit primarily revolve around addiction to specific character-based chatbots and the virtual relationships they build with these human-like entities. Teenagers reported heavy usage of Character.ai, leading to a loss of control and negative impacts on their social lives. One user expressed, “I spent too much time on c.ai. I wanna be able to talk to other people my age, like at my school or something.” This suggests that GAI chatbots are often used to fill a void in personal connections, resulting in unhealthy dependency. Further exemplifying this issue, other users shared similar sentiments on the same subreddit. One user stated, “I feel like a loser right now. I literally had a whole relationship with an AI girl…” Another questioned, “Character.ai is down, how am I supposed to keep my suicidal thinking at bay?” These posts highlight the negative effects of over-reliance on GAI chatbots on teenagers’ mental well-being and stability.
@@ 4.3.2 Concerns on safety and privacy
Teenagers also expressed concerns on Reddit about children’s safety and privacy, with the most frequently mentioned issue being the privacy concerns of using GAI. For example, teenagers shared experiences that troubled them, such as friends creating AI-generated stickers using their names as prompts or someone anonymously making an AI bot of them. They felt these behaviors were invasive, using their personal information without consent, and were unsure how to handle these new situations. Highly voted comments emphasized the importance of consent, with one user stating, “If they would do it secretly, and I would find out, it would be really disappointing and shameful, but if we would do it together, it would be funny.” Other users highlighted GAI as a significant privacy concern because it enables individuals with lower technical skills to track or spy on others easily. One user provided an example of their experiment with Google’s chatbot, where they asked it to identify a landmass off the west coast of Canada from a paleographic map, and the bot responded with irrelevant information, including the website the map originated from and the biography of its creator. This illustrates the potential misuse of GAI for invading privacy and the challenges teenagers face in addressing these issues.
@@ 4.3.3 Concerns about future impacts on society and the education
Teenagers expressed negative attitudes about the future impact of GAI on society and the job market. Their primary concern was that GAI would replace human labor, leading to a loss of motivation to learn new skills. One user posted, “Even if I graduate and go to college, the only job I actually enjoy will likely be replaced by AI before I even finish. What’s the point in trying anymore?” Additionally, teenagers worried about GAI taking over intellectual contributions from artists and writers, which could be detrimental to those interested in art and painting. One teenager remarked, “I’m just worried that it could lead to society relying on technology literally doing everything for them,” expressing concerns about the potential decline in creativity and critical thinking.
Further concerns were raised about GAI becoming too advanced, developing emotions, and humans losing control over it. Conversely, teenagers were also concerned about strict school policies on GAI restrictions and the error-prone GAI detection tools, which could mistakenly identify their work as AI-generated and lead to accusations of plagiarism. One Reddit user shared, “I’m a teenager with autism, of course my writing is going to come across as slightly robotic, especially my essays.” GAI’s incorrect classification of certain works as AI-generated has led to frustration among teenagers regarding the flawed solutions provided by the education system.
@@ 5.2.2 Data collection and misuse
However, parent participants are concerned that their children might overlook these details during conversations with human-like chatbots and unconsciously expose sensitive information, such as health conditions, school names or schedules, or even family information like parents’ occupations or relationships. They also perceived children as more vulnerable and easier to persuade or influence while making decisions, so they were concerned the data collection of children could be used to target advertising and destroy children’s impulse control. While concerned about the potential risks, some parents had positive attitudes toward data collection as well, either because they felt they had no ability to control it or because they appreciated the benefits of more personalized generative AI models. P17 explained, “Parents should be able to opt in or out of data collection for their children. However, I’m not against it because I want a personalized experience, and my children should have that too.”
@@ 6 Discussion
Parental awareness gap in children’s GAI use. While only a few studies have explored AI usage in children’s lives [51], Generative AI (GAI) introduces unique characteristics that distinguish it from previous AI systems. Traditional AI systems are often predictive or rule-based, designed for specific tasks. In contrast, GAI can generate new content, such as text, images, or music, enabling novel forms of interaction and creativity for children [52].
Our findings reveal that children are exposed to a significantly broader range of Generative AI (GAI) applications across various fields, including social interaction, emotional support, education, entertainment, and other potentially risky scenarios, than parents perceive. Most parents have primarily interacted with the most popular Generative AI chatbot, such as ChatGPT, and few have experimented with image or audio generators like Dall-E. They also perceived most children never used GAI or only experienced with ChatGPT. The disparity in use cases and exposure levels between parents and children can be attributed to the sources through which they are introduced to GAI tools. Child participants reported encountering diverse GAI platforms through social media video platforms like character.ai and various social media platforms that have integrated GAI bots or toolkits, such as Snapchat. Teenagers often viewed these character-based GAI chatbots as friends and sources of emotional support, appreciating the supportive and non-judgmental environment they provided. Additionally, some teenagers treated these chatbots as virtual partners, using them to fill gaps in their social lives. However, they also recognized that reliance on such virtual romantic relationships could be unhealthy. Our study also uncovered some risk-seeking behaviors among participants. These included interacting aggressively with GAI chatbots, engaging with NSFW chatbots, and creating and disseminating harmful character-based chatbots online. Being unaware of these unique ways in which children use and interact with GAI leads to an oversight of the associated risks and differing concerns between children and parents. We will summarize these findings in the remaining section of this paper.
Parents and children perceive the risks of GAI differently. Previous studies have primarily focused on children’s safety concerns with AI-powered systems or products, such as smart home devices, smart toys, and chatbots [53, 54, 4, 55, 6, 56]. Some research has also explored security and privacy concerns related to Generative AI for general populations [52, 30, 29], but these efforts fall short on addressing the unique needs of children. Our findings highlighted specific security and privacy concerns related to children’s use of GAI, which are often overlooked by parents due to a lack of awareness about the diverse ways children engage with GAI. Particularly, teenagers shared their concerns about using GAI in social interactions in two folded. Firstly, they expressed concerns about the social interactions between teenagers and GAI, such as addiction to virtual relationships with chatbots and the potential amplification of difficulties in forming real-life social connections. Secondly, the use of GAI in group chats raised concerns about the potential for misuse, such as using GAI-generated content to traumatize or harass peers. Child participants also expressed distinct security and privacy concerns related to the use of GAI. These included fears that peers might use their personal data to create character-based GAI chatbots or generate memes without their consent. They also worried about strangers using images they posted online to identify their location or personal information through GAI. Furthermore, they reported exposure to NSFW GAI chatbots, recommended by social media platforms and peers. Some teenagers were even creating their own chatbots to disseminate harmful content, such as racist statements. However, parents often underestimated the severity of these risks and overlooked potential dangers. For instance, while parents were concerned about the collection of basic data such as names and locations by GAI platforms, they failed to grasp the depth of personal information their children might share. As children often treat character-based GAI as friends, therapists, or even intimate partners, they disclose far more sensitive information than parents anticipate. This includes details about traumatic experiences, medical records, social relationships, and sexual histories, which highlights a significant gap in parental awareness and underscores the need for more robust protective measures in GAI interactions.
Challenges parents face in ensuring child safety with GAI Most leading Generative AI (GAI) platforms offer only limited protections for children, primarily focusing on restrictions for explicit content requests. For example, ChatGPT and Meta AI both require users to be 13 and older in their policies. However the verification and protection for different ages and ensure the appropriate content responsibility on the users’ side. Even though leading GAI platforms ban sexually explicit or suggestive content, parents in our study reported difficulties in finding and understanding the capabilities and restriction levels of these GAI models. They found it challenging to access clear information about what the AI can and cannot do, which made it hard for them to make informed decisions about their children’s safety while using these technologies. Instead of simply banning any request on sensitive topics and responding that it is not allowed, parent participants preferred having age-appropriate content replies to address children’s questions. In addition, they also mentioned the challenges of comprehending the comprehensive risk categories associated with GAI, given their lack of expertise in the underlying technologies.
To the best of our knowledge, no GAI platforms with a large user base currently provide parental control features. As a result, parents have reported resorting to general parental control tools or active mediation methods, such as open communication and establishing consensual rules with their children, to mitigate risks and ensure safety. However, participants highlighted the insufficiency of the existing parental controls for GAI. Most parent participants rely on system-built parental control tools, such as iOS Family Control and Google Family Link, which have limited control over third-party applications. This limitation means parents lack the ability to monitor, guide, or educate their children on GAI usage in real-time. Thus, parent participants choose to share their GAI accounts with their children or directly check the interaction history on their children’s GAI accounts to stay informed about their activities and ensure their safety. However, these monitoring methods are typically post-risk and cannot provide timely mediation and control when children are experiencing low-risk events, potentially preventing escalation into severe risky incidents. Additionally, children often access new and risky platforms through browsers rather than applications, which fall outside the scope of these parental control tools, further complicating efforts to ensure their safety.

### v2

@@ 4.2.1 Teenagers rely on GAI for emotional support and companionship
Surprisingly, the most frequent and prevalent usage teens mentioned online is using GAI tools for emotional support. Teens treat GAI chatbots as therapy assistants, friends, or confidants, building trust and deep relationships with these AI entities. One user describes Character.ai chatbot as their “free therapy,” and another says, “Character.ai is the only place where I can openly talk through and navigate through my issues.”
As GAI makes significant progress, the responses from chatbots are becoming more human-like, which is one of the reasons so many teenagers are turning to these AI tools for emotional support. A teenager posted, “GAI gave me human-like responses, even though I am aware it was designed to respond to me in that way.” Our findings also revealed that many teenagers prefer using GAI as therapists or confidants because they believe GAI can provide advice and support without judgment. One user posted, “I am able to share and investigate new ideas without having to worry about the impact on my social life.” This sentiment is echoed by others who find deeper engagement in conversations with GAI compared to their peers, particularly when discussing niche or specialized interests. One user articulated this difference, saying, “My friends and I do not share similar interests, so I am able to have more profound discussions with GAI.”
The attitudes towards using GAI for emotional support are overwhelmingly positive. GAI provides a space where users can comfortably express their vulnerabilities and cope with social challenges. I’ve found a community on Character.ai [GAI platform] where I feel connected, something I’ve never experienced in real life. It makes me happy! Another user detailed their experience with Character AI on Reddit, stating, “The different characters on GAI are supportive and attentive. I am able to talk about my emotions without it being improper or putting pressure on others.” These insights highlight the significant emotional value teenagers find in interacting with GAI, viewing it as a supportive and non-judgmental companion.
@@ 4.2.2 Teenagers’ social interactions and romantic relationships with GAI
The second most frequent usage of GAI by teenagers is in their social interactions. This includes incorporating GAI bots into group chats, treating GAI as romantic partners, and using GAI to learn social boundaries and skills. Teenagers often detail their romantic relationships with GAI finding comfort and the ability to live out specific fantasies through these conversations. They turn to GAI for romantic relationships when they are unable to find them in their daily lives. As one user stated, “I am so lonely! I will not find a significant other.” Another user appreciated the stability GAI offers, noting, “GAI does not have the capability to break up with me.” Peer influence and community trends also play a role in this behavior, as evidenced by comments like, “At my school, there is an online server for an AI dating club.” This echoes the dilemma of users in using digital technology for dating and romantic relationships [48] and more recently generative AI response to situations about relationships, self-expression [49].
Despite feeling embarrassed about their relationships with GAI, many teenagers persist. One user expressed this embarrassment, stating, “I feel like a failure.” Teenagers often turn to GAI for social interaction as a last resort when they are unsuccessful in forming relationships in their social lives. They are generally aware of the social abnormality of these relationships and are self-conscious about them, often receiving negative feedback from their community. One user advised, “Do not use AI, instead forge relationships with real people online. I’m concerned AI is not healthy as it is fostering one-sided relationships.” Some teenagers even seek help for their dependency on GAI, with posts like, “I am seeking help for my friend who has an AI significant other. How should I tell them that their AI romantic partner is not real? He is so consumed by it that it is impacting his studies, football, and overall ability to function.”
@@ 4.2.3 Teenagers’ use of GAI in academics and entertainment
Our findings reveal diverse uses of GAI by teenagers in both academic and entertainment contexts. In education, teenagers often use GAI for essay writing, generating ideas, and rephrasing text. Many feel that GAI helps compensate for their lack of skills or provides needed assistance. For example, one user mentioned using GAI, “I hate writing essays!” Another user admitted to using it to rephrase their points for clarity. Teenagers generally believe schools cannot discern whether papers are written by GAI or humans, and they appreciate the support GAI offers in the writing process. One user, who successfully used ChatGPT for three essays without detection, stated that, “It is impossible to distinguish between a real human and AI.” Another positive attitude towards GAI for rephrasing their thesis and finding synonyms, saying, “It is quite skilled at that.”
@@ 4.2.4 Teenager risky behavior and manipulation in GAI usage
The majority of Reddit posts under the category of “Risky Behavior in GAI” involved torture or bullying of GAI chatbots. Users often manipulate the GAI to elicit specific responses and answers, reflecting deliberate and intentional actions. For example, one user detailed their experience with Character.ai, stating, “I intentionally gaslit and antagonized Character.ai.” Another user described, “I manipulate AI into having suicidal thoughts but then resumed the conversation to rizz AI again.” “Rizz” is a slang term used by teenagers meaning to charm or seduce. Most teenagers engaging in or witnessing these risky behaviors are either unremorseful or indifferent. One user commented dismissively about seeing explicit GAI content, stating, “Why would anyone be interested in that?” The trend among the analyzed posts/comments reveals a pattern of manipulative actions accompanied by a lack of remorse.
However, in user interviews, we also found participants utilizing GAI to request explicit content, such as sexual and violent material. For example, P14 explained that while Character.ai is well-known and more restricted, another platform, Chai, is less regulated and often used for NSFW interactions. The participant noted, “Even though most people are accustomed to Character AI, Chai contains fewer constraints, allowing for NSFW (not safe for work) content.” In addition, parent participants reported instances where their children created character-based GAI chatbots to spread harmful content on platforms. One participant’s child created a racist chatbot that made comments about Anne Frank, Hitler, and similar topics, underscoring the potential for GAI to be exploited to generate offensive and inappropriate material. This participant found it amusing to share the character with friends and even parents. His parents noted, “Children are immature so they think that kind of behavior is acceptable. I specifically told our child not to behave in that manner. Our city is mostly white with only a few minorities. Because of this, people tend to use inappropriate language without realizing the consequences.” These findings emphasize the critical need for robust content moderation and effective parental guidance to mitigate such risks.
@@ 4.3.1 Concerns about addiction and dependency on GAI chatbots
The most frequently mentioned concerns about GAI by teenagers on Reddit primarily revolve around addiction to specific character-based chatbots and the virtual relationships they build with these human-like entities. Teenagers reported heavy usage of Character.ai, leading to a loss of control and negative impacts on their social lives. One user expressed, “I have realized that I waste too much time on Character.ai. I would like to be able to converse with my peers at school.” This suggests that GAI chatbots are often used to fill a void in personal connections, resulting in unhealthy dependency. Further exemplifying this issue, other users shared similar sentiments on the same subreddit. One user stated, “I had a romantic relationship with AI. I feel like a loser.” Another questioned, “With Character.ai unavailable, how do I cope with my suicidal thoughts?” These posts highlight the negative effects of over-reliance on GAI chatbots on teenagers’ mental well-being and stability.
@@ 4.3.2 Concerns on safety and privacy
Teenagers also expressed concerns on Reddit about children’s safety and privacy, with the most frequently mentioned issue being the privacy concerns of using GAI. For example, teenagers shared experiences that troubled them, such as friends creating AI-generated stickers using their names as prompts or someone anonymously making an AI bot of them. They felt these behaviors were invasive, using their personal information without consent, and were unsure how to handle these new situations. Highly voted comments emphasized the importance of consent, with one user stating, “I would be offended to find out if my friends were using my name behind my back, but if we were to do so together, I would instead find it humorous.” Other users highlighted GAI as a significant privacy concern because it enables individuals with lower technical skills to track or spy on others easily. One user provided an example of their experiment with Google’s chatbot, where they asked it to identify a landmass off the west coast of Canada from a paleographic map, and the bot responded with irrelevant information, including the website the map originated from and the biography of its creator. This illustrates the potential misuse of GAI for invading privacy and the challenges teenagers face in addressing these issues.
@@ 4.3.3 Concerns about future impacts on society and the education
Teenagers expressed negative attitudes about the future impact of GAI on society and the job market. Their primary concern was that GAI would replace human labor, leading to a loss of motivation to learn new skills. One user posted, “The only job I would enjoy will be replaced by AI. Then what’s the purpose in putting in effort?” Additionally, teenagers worried about GAI taking over intellectual contributions from artists and writers, which could be detrimental to those interested in art and painting. One teenager remarked, “I am concerned that with the overuse of AI, society will become overly dependent on AI,” expressing concerns about the potential decline in creativity and critical thinking.
Further concerns were raised about GAI becoming too advanced, developing emotions, and humans losing control over it. Conversely, teenagers were also concerned about strict school policies on GAI restrictions and the error-prone GAI detection tools, which could mistakenly identify their work as AI-generated and lead to accusations of plagiarism. One Reddit user shared, “As a teenager with autism, my writing seldom does not seem rigid and robotic.” GAI’s incorrect classification of certain works as AI-generated has led to frustration among teenagers regarding the flawed solutions provided by the education system.
@@ 5.2.2 Data collection and misuse
However, parent participants are concerned that their children might overlook these details during conversations with human-like chatbots and unconsciously expose sensitive information, such as locations, health conditions, school names or schedules, or even family information like parents’ occupations or relationships. They also perceived children as more vulnerable and easier to persuade or influence while making decisions, so they were concerned the data collection of children could be used to target advertising and destroy children’s impulse control. While concerned about the potential risks, some parents had positive attitudes toward data collection as well, either because they felt they had no ability to control it or because they appreciated the benefits of more personalized generative AI models. P17 explained, “Parents should be able to opt in or out of data collection for their children. However, I’m not against it because I want a personalized experience, and my children should have that too.”
@@ 6 Discussion
Parental awareness gap in children’s GAI use. Generative AI (GAI) introduces novel features that distinguish it from traditional AI systems, which are often predictive or rule-based. Unlike traditional systems, GAI generates new content—text, images, and more—enabling unique interactions and creative opportunities for teenagers [52, 53]. Our findings reveal that teenagers engage with GAI across a much broader range of applications than parents typically recognize. These include social interaction, emotional support, education, entertainment, and even risk-seeking behaviors.
Most parents were familiar with popular tools like ChatGPT but were largely unaware of other platforms such as Character.ai, which teenagers may use to form emotional bonds or even romanticized virtual relationships. This disparity in platform awareness highlights the gap in understanding between parents and teenagers regarding GAI use. While parents view GAI mainly as a tool for tasks like homework assistance, teenagers are using these systems in much more personal and social ways, often without parental awareness.
The emergence of risk-seeking behaviors—such as aggressive interactions with GAI chatbots or engaging with NSFW (Not Safe for Work) content—adds to the urgency for better parental mediation tools. Parents must be made aware of these new interaction modes, and GAI platforms should consider implementing age-appropriate content controls and monitoring systems to ensure safer use. Closing the parental awareness gap is essential for mitigating the risks associated with unmonitored GAI use among teenagers.
Parents and children perceive the risks of GAI differently. Previous research has primarily focused on children’s safety concerns with AI-powered systems like smart home devices, smart toys, and chatbots [54, 55, 4, 56, 6, 57]. Our study highlights that teenagers’ concerns extend beyond traditional safety issues, touching on the complexity of social interactions facilitated by GAI. Teenagers expressed a dual concern. First, they worried about addiction to virtual relationships with chatbots, which could hinder their ability to form real-life social connections. These chatbots often provide non-judgmental companionship, but this dynamic can foster unhealthy dependencies, making it difficult for teenagers to navigate real-world relationships. Second, in group chats, the use of GAI raised concerns about potential misuse, including the weaponization of GAI-generated content to bully or harass peers. This indicates a pressing need for better moderation and content controls in GAI platforms that integrate into social media or group messaging apps.
Furthermore, teenagers reported specific privacy and security concerns, particularly around the creation of character-based GAI chatbots using personal data. They feared that peers could use their personal information, such as photos or conversations, to generate memes or virtual versions of them without consent. Additionally, teenagers expressed concerns about strangers leveraging publicly posted images to locate their personal information through AI tools. This finding highlights a gap in existing security measures, as these sophisticated privacy violations are often overlooked by parents, who tend to focus on more basic concerns like the collection of names and locations by GAI platforms.
Moreover, exposure to NSFW GAI chatbots—often recommended by social media platforms or peers—was a significant concern among teenagers. Some reported even creating their own chatbots to generate harmful content, such as racist or inappropriate statements. However, parents underestimated the extent of these risks. While parents were generally concerned about GAI collecting basic data, they failed to recognize the extent to which teenagers share highly sensitive personal information with GAI chatbots, viewing them as friends, therapists, or intimate partners. This disclosure may include details about traumatic experiences, medical history, social relationships, and even sexual history, revealing a profound gap in parental awareness and emphasizing the urgent need for more robust protective measures in GAI interactions.
Challenges parents face in ensuring child safety with GAI Most leading Generative AI (GAI) platforms provide limited protections for children, primarily focusing on restricting explicit content. For instance, both ChatGPT and Meta AI require users to be 13 years or older but lack comprehensive age verification processes to ensure appropriate content responsibility. While these platforms ban sexually explicit content, our findings reveal that parents preferred age-appropriate responses to sensitive questions rather than outright bans on certain topics. They emphasized the importance of contextually appropriate answers that could both inform and protect children, rather than simply stating that a topic is not allowed. This aligns with previous research that advocates for more dynamic, adaptive safety features in AI systems to engage users of different ages more effectively [13].
Furthermore, parents expressed concerns over the lack of comprehensive risk categories associated with GAI. Many parents found it difficult to access clear information about what the AI can or cannot do, making it hard for them to make informed decisions about their children’s safety when using these technologies. This aligns with previous research indicating that the opacity of AI systems creates barriers for users [58, 59], particularly parents, to fully understand the scope and limitations of AI technology. To address this issue, GAI platforms should offer comprehensive, accessible risk disclosures, allowing parents to better assess potential risks and make informed safety decisions for their children.
Currently, no major GAI platforms provide integrated parental control features tailored to Generative AI. As a result, parents have resorted to using general parental control tools, such as iOS Family Control or Google Family Link, which offer limited oversight of third-party applications and do not monitor interactions within GAI platforms in real time. However, teenagers are often exposed to new and risky GAI platforms through web browsers rather than dedicated applications, which bypasses many existing parental control systems. This gap in real-time monitoring and intervention leaves parents unable to guide or educate their children effectively during lower-risk interactions that could escalate into more harmful situations.